When each side of a square in increased by 2 cm., the area is increased by 28 sq. cm. What is the perimeter of the original square?
Answer by ankor@dixie-net.com(22740)

When each side of a square in increased by 2 cm., the area is increased by 28 sq. cm.
What is the perimeter of the original square?
:
let x = the side of the square
(x+2)^2 = x^2 + 28
FOIL(x+2)(x+2)
x^2 + 4x + 4 = x^2 + 28
Subtract x^2 from both sides
4x + 4 = 28
4x = 28 - 4
4x = 24, this would be the perimeter
x = 24/4
x = 6 cm on each side
" What is the perimeter of the original square?"
4(6) = 24 cm